### v2

@@ 5.1 Theoretical Foundation
Although NE has been shown to exist in games with finite players and finite actions [100], computing NE with self-play algorithms in larger games remains challenging and consequently, many studies aim to achieve approximate NE [76]. However, in some cases, even computing an approximate NE is difficult [31]. Some research has resorted to higher levels of equilibrium, such as CE [86] and $\alpha$-rank [96]. Although many algorithms are developed with theoretical foundations, there is often a gap in complex real-world scenarios. For example, although self-play algorithms like AlphaGo [115], AlphaStar [109], and OpenAI Five [7] achieve empirical success, they lack formal game theory proofs behind their effectiveness. Future work should focus on bridging this gap by developing new theoretically sound and practically effective algorithms or proving the theoretical underpinnings of existing successful self-play algorithms in complex environments.
@@ 5.2 Non-stationarity of the Environment
In the self-play framework, the opponents are a vital component of the environment, and the strategies of the opponent players evolve as training progresses. This evolution can cause the same strategy to lead to different results over time, creating a non-stationary environment. This problem is also shared by the MARL area. Future research should aim to develop self-play algorithms that are more robust and can adapt to changing conditions. For example, incorporating opponent modeling [150] into self-play can help agents anticipate changes in opponent strategies and adjust their own strategies proactively, making them more robust to environmental changes.
@@ 5.3 Scalability and Training Efficiency
The scalability of self-play methods faces significant challenges as the number of teams and players within those teams increases. As the number of participants grows, the complexity of interactions explodes. For example, in OpenAI Five [7], the hero pool size is limited to only 17 heroes. MOBA AI [143] extends this to a 40-hero pool with the help of curriculum learning, but it still cannot cover the entire hero pool available in the actual game. One potential solution is leveraging players’ inherent connections to optimize the learning process. For instance, using graph-based models to represent and exploit the relationships between players can help manage and reduce the complexity of large-scale multi-agent environments. These scalability issues are fundamentally rooted in the limited training efficiency of self-play methods from two aspects including computation and storage. The first issue is computational efficiency, which is induced by the iterative nature of self-play where agents repeatedly play against themselves or past versions. Moreover, although forming more complex populations and competitive mechanisms [109] can enhance the intensity and quality of training, it further exacerbates the demand for computational resources. Techniques such as parallel computing, distributed learning, and more efficient neural network architectures could be explored to address these challenges. The second issue is storage because self-play requires maintaining a policy pool. Even when using a shared network architecture, storing the parameters of large models can be problematic. This issue is particularly pronounced in regret-minimization-based self-play algorithms, which must store the regrets for each information set and potential action. Managing both the computational load and storage requirements is essential for improving the overall training efficiency and scalability of self-play methods.
@@ 5.4 With Large Language Models
With their remarkable capability and emergent generalizability, large language models (LLMs) have been regarded as a potential foundation for achieving human-level intelligence [1], and self-play methods have been proposed to fine-tune LLMs, enhance LLMs’ reasoning performance, and build LLM-based agents with strong decision-making abilities. Post-training fine-tuning [102, 3] is a key step in aligning LLM with more desired behaviors, but it requires a huge amount of human-annotated data. To reduce reliance on human-annotated data, Self-play fIne-tuNing (SPIN) [26] introduces a self-play mechanism to generate training data using the LLM itself and fine-tuning the LLM by distinguishing self-generated responses from human-annotated data. Another line of work [98, 121, 135] formulate the alignment problem as a two-player constant-sum game and use self-play methods to solve the game. Some other work [56, 146] also utilizes model-generated data to fine-tune LLMs with minimal human annotations. The idea of self-improvement has also been applied to improve the reasoning ability of LLMs. Self-Play of Adversarial Game (SPAG) [27] observes that self-play in a two-player adversarial language game called Adversarial Taboo can boost the LLM’s performance on various reasoning benchmarks. Besides improving the capability of LLMs, self-play methods have also contributed to building LLM-based agents with strong strategic abilities. A representative work is Cicero [38] which achieves human-level play in the game of Diplomacy by combining language models with an RL policy trained by self-play. Cicero uses the self-play policy to produce an intended action and prompts the language model to generate languages conditioned on the policy’s intent. Another work [138] also combines LLM with self-play policy but takes a different approach by first prompting the LLM to propose multiple action candidates and then using the self-play policy to produce the action distribution over these candidates. Despite recent progress, applying self-play with LLMs is still underexplored and requires further research.
@@ 5.5 Real-World Applications
Self-play is a powerful technique with widespread applications across various domains. It is particularly effective in addressing some problems abstracted from real-world situations through its iterative learning approach. In the field of economics, for instance, self-play is used to enhance supervised learning models in multi-issue bargaining tasks [74]. Furthermore, self-play proves beneficial in solving combinatorial optimization problems (COPs) such as the traveling salesman problem (TSP) and the capacitated vehicle routing problem (CVRP) [130]. Within the domain of traffic, self-play facilitates the development of human-like autonomous driving behaviors [28] and enables vehicles to learn negotiation strategies to merge on or off roads [123], although currently within 2D simulators.
However, a notable challenge with self-play is its impracticality for direct application in real-world scenarios. Its iterative approach requires extensive trial and error, which is computationally demanding and often involves actions that are impractical or unsafe outside of a controlled simulation. Therefore, self-play is often carried out in simulators. The effective deployment of self-play in real-world applications boils down to overcoming the Sim2Real gap. For example, for tasks where the Sim2Real gap is less pronounced, self-play can be well employed to enhance video streaming capabilities [59], and to address the image retargeting problem [66]. Moreover, EvoPlay [132] leverages self-play to design protein sequences, utilizing the advanced AlphaFold2 simulator [65] to narrow the Sim2Real gap. Similarly, in heterogeneous multi-robot systems, self-play is utilized for competitive sports tasks like humanoid football [81, 47], quadruped competition [136], and multi-drone volleyball [139], with significant efforts dedicated to Sim2Real transitions for real-world success [41].
@@ 6 Conclusion
Acknowledgment
This research was supported by National Natural Science Foundation of China (No.62406159, 62325405), Postdoctoral Fellowship Program of CPSF under Grant Number (GZC20240830, 2024M761676), China Postdoctoral Science Special Foundation 2024T170496.

### v3

@@ 5.1 Theoretical Foundation
Although NE has been shown to exist in games with finite players and finite actions (Nash et al., 1950), computing NE with self-play algorithms in larger games remains challenging and consequently, many studies aim to achieve approximate NE (Li et al., 2024). However, in some cases, even computing an approximate NE is difficult (Daskalakis, 2013). Some research has resorted to higher levels of equilibrium, such as CE (Marris et al., 2021) and $\alpha$-rank (Muller et al., 2020). Although many algorithms are developed with theoretical foundations, there is often a gap in complex real-world scenarios. For example, although self-play algorithms like AlphaGo (Silver et al., 2016), AlphaStar (Samvelyan et al., 2019), and OpenAI Five (Berner et al., 2019) achieve empirical success, they lack formal game theory proofs behind their effectiveness. Future work should focus on bridging this gap by developing new theoretically sound and practically effective algorithms or proving the theoretical underpinnings of existing successful self-play algorithms in complex environments.
@@ 5.2 Non-stationarity of the Environment
In the self-play framework, the opponents are a vital component of the environment, and the strategies of the opponent players evolve as training progresses. This evolution can cause the same strategy to lead to different results over time, creating a non-stationary environment. This problem is also shared by the MARL area. Future research should aim to develop self-play algorithms that are more robust and can adapt to changing conditions. For example, incorporating opponent modeling (Zhao et al., 2022b) into self-play can help agents anticipate changes in opponent strategies and adjust their own strategies proactively, making them more robust to environmental changes.
@@ 5.3 Scalability and Training Efficiency
The scalability of self-play methods faces significant challenges as the number of teams and players within those teams increases. As the number of participants grows, the complexity of interactions explodes. For example, in OpenAI Five (Berner et al., 2019), the hero pool size is limited to only 17 heroes. MOBA AI (Ye et al., 2020a) extends this to a 40-hero pool with the help of curriculum learning, but it still cannot cover the entire hero pool available in the actual game. One potential solution is leveraging players’ inherent connections to optimize the learning process. For instance, using graph-based models to represent and exploit the relationships between players can help manage and reduce the complexity of large-scale multi-agent environments. These scalability issues are fundamentally rooted in the limited training efficiency of self-play methods from two aspects including computation and storage. The first issue is computational efficiency, which is induced by the iterative nature of self-play where agents repeatedly play against themselves or past versions. Moreover, although forming more complex populations and competitive mechanisms (Samvelyan et al., 2019) can enhance the intensity and quality of training, it further exacerbates the demand for computational resources. Techniques such as parallel computing, distributed learning, and more efficient neural network architectures could be explored to address these challenges. The second issue is storage because self-play requires maintaining a policy pool. Even when using a shared network architecture, storing the parameters of large models can be problematic. This issue is particularly pronounced in regret-minimization-based self-play algorithms, which must store the regrets for each information set and potential action. Managing both the computational load and storage requirements is essential for improving the overall training efficiency and scalability of self-play methods.
@@ 5.4 With Large Language Models
With their remarkable capability and emergent generalizability, large language models (LLMs) have been regarded as a potential foundation for achieving human-level intelligence (Achiam et al., 2023), and self-play methods have been proposed to fine-tune LLMs, enhance LLMs’ reasoning performance, and build LLM-based agents with strong decision-making abilities. Post-training fine-tuning (Ouyang et al., 2022; Bai et al., 2022) is a key step in aligning LLM with more desired behaviors, but it requires a huge amount of human-annotated data. To reduce reliance on human-annotated data, Self-play fIne-tuNing (SPIN) (Chen et al., 2024) introduces a self-play mechanism to generate training data using the LLM itself and fine-tuning the LLM by distinguishing self-generated responses from human-annotated data. Another line of work (Munos et al., 2024; Swamy et al., 2024; Wu et al., 2024) formulate the alignment problem as a two-player constant-sum game and use self-play methods to solve the game. Some other work (Huang et al., 2022; Yuan et al., 2024) also utilizes model-generated data to fine-tune LLMs with minimal human annotations. The idea of self-improvement has also been applied to improve the reasoning ability of LLMs. Self-Play of Adversarial Game (SPAG) (Cheng et al., 2025) observes that self-play in a two-player adversarial language game called Adversarial Taboo can boost the LLM’s performance on various reasoning benchmarks. Besides improving the capability of LLMs, self-play methods have also contributed to building LLM-based agents with strong strategic abilities. A representative work is Cicero († et al.(2022)(FAIR)†, Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu, et al., FAIR) which achieves human-level play in the game of Diplomacy by combining language models with an RL policy trained by self-play. Cicero uses the self-play policy to produce an intended action and prompts the language model to generate languages conditioned on the policy’s intent. Another work (Xu et al., 2024) also combines LLM with self-play policy but takes a different approach by first prompting the LLM to propose multiple action candidates and then using the self-play policy to produce the action distribution over these candidates. Despite recent progress, applying self-play with LLMs is still underexplored and requires further research.
@@ 5.5 Real-World Applications
Self-play is particularly effective in addressing some problems abstracted from real-world situations through its iterative learning approach. In the field of economics, for instance, self-play is used to enhance supervised learning models in multi-issue bargaining tasks (Lewis et al., 2017). Furthermore, self-play proves beneficial in solving combinatorial optimization problems (COPs) such as the traveling salesman problem (TSP) and the capacitated vehicle routing problem (CVRP) (Wang et al., 2024). Within the domain of traffic, self-play facilitates the development of human-like autonomous driving behaviors (Cornelisse & Vinitsky, 2024) and enables vehicles to learn negotiation strategies to merge on or off roads (Tang, 2019), although currently within 2D simulators.
However, a notable challenge with self-play is its impracticality for direct application in real-world scenarios. Its iterative approach requires extensive trial and error, which is computationally demanding and often involves actions that are impractical or unsafe outside of a controlled simulation. Therefore, self-play is often carried out in simulators. The effective deployment of self-play in real-world applications boils down to overcoming the Sim2Real gap. For example, for tasks where the Sim2Real gap is less pronounced, self-play can be well employed to enhance video streaming capabilities (Huang et al., 2021b), and to address the image retargeting problem (Kajiura et al., 2020). Moreover, EvoPlay (Wang et al., 2023) leverages self-play to design protein sequences, utilizing the advanced AlphaFold2 simulator (Jumper et al., 2021) to narrow the Sim2Real gap. Similarly, in heterogeneous multi-robot systems, self-play is utilized for competitive sports tasks like humanoid football (Liu et al., 2022b; Haarnoja et al., 2024), quadruped competition (Xiong et al., 2024), and multi-drone volleyball (Xu et al., 2025), with significant efforts dedicated to Sim2Real transitions for real-world success (Gao et al., 2023).